OpenStudy (anonymous):

a rectangle has a length of x+5 and a with of 4x-6. Which expression describes the area of the rectangle a. 5x-1 b. 4x^2-1 c. 4x^2+14x+30 d. 4x^2+14x-30

OpenStudy (opcode):

A rectangle's area is \[= w \times h\] So: \[(x + 5)(4x - 6)\] FOIL that and you should have your answer.

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

4x^2+14x-30

OpenStudy (whpalmer4):

not that you should use it here, but a quick shortcut to multiplying the whole thing out is to remember that when multiplying polynomials, the constant term (the plain number) will be the product of all of the constant terms in the product terms. \[(x+3)(x-5) = x*x-5x+3x-15 = x^2-2x-15\]If you had a bunch of choices, all with different constant terms, 3*(-5) = -15, pick the one that matches :-)